# A Survey on Brain-inspired Deep Learning via Predictive Coding

### v2

@@ 3 Popular Implementations of Predictive Coding
Informally, PC claims that there exist two families of neurons that define an internal (i.e., generative) model of the world: the first generates predictions that are passed to lower layers, and the second encodes prediction errors that are passed to higher layers [20, 46]. To discuss this basic PC architecture, in this section, we define three quantities, referred to as the value node, the error node, and the prediction of every neuron. The value node $\mathbf{x}^{\ell}_{i,t}$ (where the indices $i$ and $\ell$ refer to the $i$-th neuron of the $\ell$-th layer at time $t$) encodes the most likely value of some latent state. The second computational unit is the prediction $\mathbf{u}^{\ell}_{i,t}$, which is a function of the value nodes of a higher level in the hierarchy, and is computed as follows:
@@
where $\phi$ is a nonlinear function (akin to the activation function in deep neural networks), and $\mathbf{W}^{\ell+1}$ is a matrix containing the predictive synaptic parameters for layer $\ell+1$. The third computational unit is the prediction error $\mathbf{e}^{\ell}_{i,t}$, given by the difference between its value node and its prediction node, i.e., $\mathbf{e}^{\ell}_{i,t}=\mathbf{x}^{\ell}_{i,t}-\mathbf{u}^{\ell}_{i,t}$. This local definition of error, which exists in every network neuron, foregrounds a key difference between PC and models trained with backprop (e.g., the multilayer perceptron), as it enables learning through only local computations. A graphical example depicting the differences in locality between PC and backprop is given in Fig. 3. Taken together, the above three quantities, as well as the set of synaptic weight matrices $(\mathbf{W}^{0},\dots,\mathbf{W}^{L})$, define a generative model where both inference and learning are performed as a means to minimize a single (global) energy function, formally defined as the sum of the squared prediction errors of every neuron:
$\mathcal{E}_{t}=\frac{1}{2}\sum_{i,l}(\mathbf{e}^{\ell}_{i,t})^{2}.$ (4)
This energy function is exactly the variational free energy defined in the previous section, and a proper derivation can be found in [47]. Let us assume that our generative model is presented with an observation $\mathbf{o}\in\mathbb{R}^{J_{0}}$. Then, the following process describes the credit assignment and consequent update of the synaptic parameters of the model: first, the neurons of the lowest layer are set equal to the sensory observation, i.e., $\mathbf{x}^{0}=\mathbf{o}$. Next, the unconstrained neural activities are updated until convergence (or, for a fixed number $T$ of iterations) to minimize the energy of Eq. 4 via gradient descent. In particular, the equation for the update dynamics of the value nodes is the following:
@@
where $\alpha$ is the learning rate for the synaptic update. The alternation of these two phases, i.e., the value node update and weight update steps, defines the learning algorithm used to train PC networks [20, 46]. Importantly, although every computation is local, both update rules minimize the same energy function, which is globally defined over the whole network. The difference in locality between a weight update of backprop and the one for PC is given in Fig. 3(a).
@@
0:  $\mathbf{x}^{L}$ is fixed to $\mathbf{o}$ for every time step $t$.
@@
3:        Update $\mathbf{x}^{\ell}_{i,t}$ to minimize $\mathcal{E}_{t}$ via Eq. (5
@@
In the above, we have described how to use PC to perform unsupervised learning tasks. However, this algorithm or variations of it have been shown to obtain a performance comparable to deep networks learned with backprop on supervised learning tasks [46, 48, 49]. To extend the above formulation to the domain of labeled data, we have to re-imagine the PC network as generating the label $\mathbf{y}$ and that the data point $\mathbf{o}$ serves as a prior on the value of the neurons in the first layer of the network. Let us assume that we have a labeled data point $(\mathbf{o},\mathbf{y})$, where, again, $\mathbf{y}$ is the label  [^1] In supervised learning, the total energy of the model can also be decomposed as $\mathcal{E}=\tilde{\mathcal{E}_{t}}+\mathcal{L}$, with $\tilde{\mathcal{E}_{t}}$ be the total energy of all the internal neurons only, and $\mathcal{L}$ be the squared error defined on the label, as in standard regression tasks. This allows to define interesting similarities with backprop [50].. Supervised learning is then realized by fixing the value nodes of the first and last layer to the entries of the data point and its label, respectively, i.e., $\mathbf{x}_{0,t}=\mathbf{o}$ and $\mathbf{x}_{L,t}=\mathbf{y}$ for every time step $t$. A graphical depiction of the difference between supervised and unsupervised learning using PC is given in Fig. 4.
@@
While defining this generalization on fully connected graphs makes the notation simpler, interesting use-cases can be found in graphs that are not fully connected. To this end, it suffices to note that every graph is a subset of a fully connected one and, hence, we can consider sparse weight matrices $\mathbf{W}$, where only the entries of the edges that we need to define our structure are non-zero. This can be done by multiplying $\mathbf{W}$ by an adjacency matrix. In the original work along these lines, the authors trained a network of multiple, sparsely-connected brain regions and tested the resulting model on image denoising and image completion tasks. The exact structure of this network is that of the Assembly Calculus, a Hebbian learning framework/method specifically designed to model brain regions [52]. A graphical sketch of how to derive such a model is given in Fig. 3(b).
@@
A body of literature has investigated the efficacy of PC networks in memory-related tasks. Specifically, this literature studies their capacity to store a dataset and retrieve individual data points when a related cue is presented. In practice, this cue is often an incomplete or corrupted version of a stored memory and the task is successfully completed if and only if the correct data point is retrieved. PC schemes have demonstrated their ability to effectively store and retrieve complex memories, such as images from the COIL and ImageNet datasets [53, 54, 55]. However, while their retrieval capabilities are robust, their capacity is still limited, and hence not comparable to that of continuous or universal Hopfield networks [56, 57]. This capacity can be improved by implementing a fast and powerful memory writing operation that allows these models to store individual memories without overwriting existing ones. Models that incorporate memory writing also tend to facilitate the implementation of a forget operation that erases individual memories with little to no impact on overall model performance [58]. It is also possible to store memories in the latent variable $\mathbf{x}_{L}$, by learning a Gaussian mixture prior, whose centers correspond to each of the the individual memories [59]. This has also been implemented in the temporal domain, where the goal is to retrieve missing future frames of a video, given initial ones [55] (as a sort of priming stimulus).
3.1 Theoretical Results on PC
@@
Here, we present theoretical results concerning the convergence properties of PC, as well as results that compare their learning capabilities against standard artificial neural networks trained with backprop. It is a common trend in the literature to compare the performance of biologically plausible algorithms to those of backprop, especially in supervised learning scenarios. Recent work has shown that PC can approximate the weight update of backprop, under specific conditions, on both MLPs (multilayer perceptrons) and within the more general framework of computational graphs [46, 60]. These conditions are restrictive in practice, as these results only hold if either the total prediction error on the network is infinitesimally small or the predictions are kept constant throughout the whole duration of the inference process, i.e., $\mathbf{u}^{\ell}_{i,t}=\mathbf{u}^{\ell}_{i,0}$ for every time step $t$. In practice, however, empirical studies have shown that for the approximation to hold, it suffices to have a small output error [60].
A similar line of work has also shown that simply adding a temporal scheduling to the updates of the weights and the neural activities leads to a weight update that is equivalent to that of backprop. This temporal scheduling, however, is only possible by adding an external control that triggers the updates at different time steps, hence reducing the biological plausibility of PC. Again, these results were first obtained in the setting of feedforward networks, and were later extended to the more general framework of computational graphs [61, 62]. Similar works have also investigated similarities between PC and other neuroscience-inspired algorithms, such as contrastive Hebbian learning [63], target propagation [64, 65, 66], and equilibrium propagation [15, 50].
Robustness.
@@
PC has been shown to perform better than standard models (such as backprop-trained deep neural networks) on problems that are faced by biological organisms, such as continual learning, online learning, and learning from a small amount of data [67, 68, 27]. This is due to the inference phase, which allows the error to be distributed in the network in a way that avoids a phenomenon called weight interference, where the update rules of the single parameters do not consider the updates of the other parameters, leading to less stable training progress. Formally, PC networks trained for supervised learning naturally model implicit gradient descent [28], a more stable formulation that reduces numerical instability by using implicit gradients, instead of explicit ones. Specifically, a parameter update of implicit gradient descent is defined as follows:
@@
This formulation is called implicit, since $\mathbf{W}_{t+1}$ appears on both sides of the equation, desirably reducing the sensitivity to the learning rate [69]. As a consequence, PC models tend to be more robust and better calibrated than standard ones, as has been shown in multiple classification tasks on both convolutional and graph networks [51, 70, 48].
@@ 3.2 Neural Generative Coding
Despite being developed to model information processing in different brain areas, the above formulation is still biologically implausible, the main implausibility being symmetric weight connections. In what follows, we survey a generalization of PC, to arbitrary wiring patterns, called neural generative coding (NGC). This variation has its roots grounded in cable theory [71, 72] and neuronal compartments [73]. As in PC, NGC instantiates a form of predict-then-correct learning and inference but crucially expands the iterative neural dynamics to incorporate other known neurobiological mechanisms, such as lateral competition, learned precision-weighting, and membrane potential leakage. Furthermore, NGC tackles the weight symmetry assumption inherent to standard PC by decoupling the forward generative pathway(s) (which are typically composed of matrices $\mathbf{W}^{\ell}$) from the message passing feedback pathway(s) (which are composed of matrices $\mathbf{E}^{\ell}$); this easily allows the introduction of skip connections in one or both pathways to create heterarchical structures [24]. Usefully, this means that NGC circuit models do not require that the feedback message passing connectivity structure mirror (in reverse) the same structure as the underlying generative model itself, offering a high degree of flexibility in designing the generative model and its requisite message passing pathways [24, 74, 75]. Furthermore, each error feedback matrix $\mathbf{E}^{\ell}$ can be updated with local rules.
Compactly, given clamped sensory vectors, an NGC circuit undergoes a settling cycle where it processes input data signals over a stimulus window of $T$ time steps. As in PC, each region of activity in an NGC circuit makes a local prediction of a nearby region (i.e., layer $\ell+1$ predicts $\ell$ in a typical hierarchical structure) as follows:
$\mathbf{\bar{x}}^{\ell}_{t}$ $=g^{\ell}\Big{(}\mathbf{W}^{\ell+1}\cdot\phi^{\ell+1}(\mathbf{x}^{\ell+1}_{t})+\alpha_{m}\big{(}\sum^{K}_{k=2}\mathbf{M}^{\ell+k}\cdot\phi^{\ell+k}(\mathbf{x}^{\ell+k}_{t}\big{)}\Big{)}$ (9)
$\mathbf{e}^{\ell}_{t}$ $=(\mathbf{\Sigma}^{\ell})^{-1}\cdot(\mathbf{x}^{\ell}_{t}-\mathbf{\bar{x}}^{\ell}_{t})\,,$ (10)
where $\mathbf{x}^{0}_{t}=\mathbf{o}$ (the bottom layer is set to the observation) and $g^{\ell}$ is a nonlinearity applied to a layer’s predictive outputs. $\mathbf{W}^{\ell}$ contains the generative synapses between layer $\ell+1$ and $\ell$, $\mathbf{M}^{\ell+k}$ houses skip generative synapses that connect layer $\ell+k$ to $\ell$, and $\mathbf{\Sigma}^{\ell}$ is the cross-correlational matrix meant to modulate the error unit signals at each time step and is specifically constrained to ensure that its main diagonal is non-negative. The activities of the internal neurons themselves (i.e., all the neurons in between the clamped layers $\ell=L\ldots 0$) are updated according to the following equation (formulated for a single layer $\ell$ in vector form):
$\mathbf{x}^{\ell}_{t+\Delta t}=\mathbf{x}^{\ell}_{t}+\beta\Big{(}-\gamma\mathbf{x}^{\ell}_{t}-\mathbf{e}^{\ell}_{t}+(\mathbf{E}^{\ell}\cdot\mathbf{e}^{\ell-1}_{t})\odot f_{D}(\mathbf{x}^{\ell}_{t})+\Phi(\mathbf{x}^{\ell}_{t})\Big{)}\;\mbox{where }\ell>0\,,$ (11)
where $\mathbf{E}^{\ell}$ is a matrix containing error feedback synapses that are meant to pass prediction errors from layer $\ell-1$ to $\ell$. Notice that while the above update is formulated as a single Euler integration step, one could employ other methods, e.g., the midpoint method, to calculate a new state update for time $t+\Delta t$. $\beta$ is the neural state update coefficient (typically set according to $\beta=\frac{\Delta t}{\tau}$, where $\tau$ is the state unit time constant, similar to a cellular membrane potential time constant, in the order of milliseconds), and $f_{D}(\cdot)$ is a dampening function (which offers the update scaling benefits that $\phi^{\prime}(\cdot)$ provides in standard PC but without requiring differentiable activation functions). $\Phi(.)$ is a special lateral interaction function that can take many possible forms, ranging from a kurtotic distributional prior placed over the neural activities, as was classically done in [76, 20], or in the form of explicit lateral synapses [24], as presented below:
$\Phi(\mathbf{x}^{\ell}_{i,t})=-\Big{(}\sum_{j\in J_{\ell},j\neq i}\mathbf{V}^{\ell}_{i,j}\phi^{\ell}(\mathbf{x}^{\ell}_{j,t})\Big{)}+\mathbf{V}^{\ell}_{ii}\phi^{\ell}(\mathbf{x}^{\ell}_{i,t})\,,$ (12)
where $\mathbf{V}^{\ell}$ is a matrix containing the lateral cross-inhibitory and self-excitation synapses for layer/region $\ell$. Note that the state update equation (Equation 11) indicates that a vector of neural activity values changes, at each step within a settling cycle, according to (from left to right), a leak term/variable (the strength of which is controlled by $\gamma$), a combined top-down and bottom-up pressure from mismatch signals in nearby neural regions/layers, and a lateral interaction term.
After processing the sensory input for $T$ time steps (repeatedly applying Equation 11 for $T$ times in a row), the synapses are then adjusted with a multi-factor Hebbian update:
$\Delta\mathbf{W}^{\ell}=\mathbf{e}^{\ell}_{t}\cdot(\phi^{\ell+1}(\mathbf{x}^{\ell+1}_{t}))^{T},\;\text{and},\;\Delta\mathbf{E}^{\ell}=\gamma_{e}(\phi^{\ell+1}(\mathbf{x}^{\ell+1}_{t}\cdot(\mathbf{e}^{\ell}_{t})^{T})\,,$ (13)
where $\gamma_{e}$ is a factor that controls the time-scale over which the error synapses are evolved ($\gamma_{e}$ is typically less than one to ensure that the error feedback synapses change a bit more slowly than the generative synapses). Note that after the Hebbian updates prescribed by Equation 13 are applied to their corresponding weight matrices, either a form of synaptic scaling is applied or the synaptic matrices are instead normalized after an update [24] (much as was done in classical sparse coding), i.e., $\mathbf{W}^{\ell}\leftarrow\mathbf{W}^{\ell}/\mid|\mathbf{W}^{\ell}||_{2}$ and $\mathbf{E}^{\ell}\leftarrow\mathbf{E}^{\ell}/||\mathbf{E}^{\ell}||_{2}$.
Another important functionality of an NGC circuit is the ability to project a vector through its underlying directed generative model ancestrally. In essence, this amounts to a feedforward pass, since no settling process is required. Formally, ancestrally projecting any vector $\mathbf{x}_{\epsilon}$ through an NGC circuit is done as follows:
$\mathbf{x}^{\ell}_{t}=\mathbf{\bar{x}}^{\ell}_{t}$ $=g^{\ell}(\mathbf{W}^{\ell+1}\cdot\phi^{\ell+1}(\mathbf{x}^{\ell+1}_{t})),\;\forall\ell=(L-1),\ldots,0\,,$ (14)
where $\mathbf{z}^{L}_{t}=\mathbf{x}_{\epsilon}$ and $\mathbf{x}_{\epsilon}$ is a sample from the NGC circuit’s prior $p(\mathbf{z}^{L})$ or an explicit top-down vector that is clamped to layer $L$. Note that since the prior over a latent state $p(\mathbf{z}^{L})$ in an NGC circuit is unconstrained and multimodal, it is common to retro-fit a mixture model to samples of $p(\mathbf{z}^{L})$ in order to facilitate fast downstream ancestral sampling [24].
3.3 Biased Competition and Divisive Input Modulation (BC-DIM)
In the biased competition (BC) model of PC [77, 78], the neural dynamics generally involves a skip connection where a layer above the current one further mediates the state update:
$\mathbf{x}^{\ell}_{t}$ $=\mathbf{x}^{\ell}_{t}+\beta(\mathbf{W}^{\ell})^{T}\cdot\mathbf{e}^{\ell-1}+\beta_{m}(\mathbf{V}^{\ell+1}\cdot\mathbf{x}^{\ell+1})+\beta_{a}(\mathbf{A}^{\ell}\cdot\mathbf{x}^{\ell,A})$ (15)
$\mathbf{e}^{\ell}_{t}$ $=\mathbf{x}^{\ell}-(\mathbf{W}^{\ell+1}\cdot\mathbf{x}^{\ell+1})\,,$ (16)
where we notice that, in this implementation, the error unit calculation remains the same as in standard PC and NGC. Note that $\mathbf{A}^{\ell}$ is a third (optional) synaptic matrix introduced to map an external set of attentional feature values $\mathbf{x}^{\ell,A}$ (for example, command control signals to induce an action-oriented behavior or labels to encourage the acquisition of discriminative latent activities). The coefficients $\beta_{m}$ and $\beta_{a}$ weigh the skip modulation and attentional feature pressures, respectively. Crucially, this model was later generalized/extended to incorporate an important neural mechanism known as divisive modulation (DIM) [79]:
$\mathbf{e}^{\ell}_{t}=\mathbf{x}^{\ell}_{t}\oslash\big{(}\epsilon_{1}+(\mathbf{W}^{\ell+1})\cdot\mathbf{x}^{\ell+1}_{t})\big{)}\,,$ (17)
with the resulting neural dynamics becoming:
$\mathbf{x}^{\ell}_{t}$ $\leftarrow(\epsilon_{2}+\mathbf{x}^{\ell}_{t})\odot\big{(}(\mathbf{W}^{\ell})^{T}\cdot\mathbf{e}^{\ell}\big{)}$ (18)
$\mathbf{x}^{\ell}_{t}$ $\leftarrow\mathbf{x}^{\ell}_{t}\odot\big{(}1+\beta((\mathbf{V}^{\ell+1})\cdot\mathbf{x}^{\ell+1}+(\mathbf{A}^{\ell})\cdot\mathbf{x}^{\ell,A})\big{)}\,,$ (19)
where we have (for consistency with the other PC frameworks) flipped the use of the transpose over the generative synaptic matrix $\mathbf{W}^{\ell}$, compared to the model’s original presentation in [79]. Furthermore, in BC-DIM circuits, $\mathbf{V}^{\ell}=(n(\mathbf{W}^{\ell})/)^{T}$ is the transpose of the normalized forward synaptic matrix $\mathbf{W}^{\ell}$, while $\epsilon_{1}$ and $\epsilon_{2}$ are small coefficients used to prevent division or multiplication by zero (to ensure numerical stability).

### v3

@@ 3 Implementations of Predictive Coding
Informally, PC claims that there exist two families of neurons that define an internal (i.e., generative) model of the world: the first generates predictions that are passed to lower layers, and the second encodes prediction errors that are passed to higher layers [20, 53]. Recent work has applied this basic separation to prediction and prediction errors within the compartments of a single neuron, where errors are propagated back via dendritic connections [54, 55]. In this work, we use the standard formulation of PC papers in machine learning, and describe an artificial neuron as a computational unit with three quantities: its value (value node), its prediction, and its error (error node), defined as the difference of the first two. The value node $\mathbf{x}^{\ell}_{i,t}$ (where the indices $i$ and $\ell$ refer to the $i$-th neuron of the $\ell$-th layer at time $t$) encodes the most likely value of some latent state. The second computational unit is the prediction $\mathbf{u}^{\ell}_{i,t}$, which is a function of the value nodes of a higher level in the hierarchy. Let $J_{\ell}$ denote the dimension (number of neurons) of layer $\ell$, $\mathbf{W}^{\ell+1}$ be a matrix containing the predictive synaptic parameters for layer $\ell+1$, and $\phi$ be a nonlinear activation function. Then, the prediction is computed as follows:
@@
where $\phi$ is a nonlinear function (akin to the activation function in deep neural networks). The third computational unit is the prediction error $\mathbf{e}^{\ell}_{i,t}$, given by the difference between its value node and its prediction node, i.e., $\mathbf{e}^{\ell}_{i,t}=\mathbf{x}^{\ell}_{i,t}-\mathbf{u}^{\ell}_{i,t}$. This local definition of error, which exists in every network neuron, foregrounds a key difference between PC and models trained with backprop (e.g., the multilayer perceptron), as it enables learning through only local computations. A graphical example depicting the differences in locality between PC and backprop is given in Fig. 3. Taken together, the above three quantities, as well as the set of synaptic weight matrices $(\mathbf{W}^{0},\dots,\mathbf{W}^{L})$, define a generative model where both inference and learning are performed as a means to minimize a single (global) energy function, formally defined as the sum of the squared prediction errors of every neuron:
$\mathcal{E}_{t}=\frac{1}{2}\sum_{i,\ell}(\mathbf{e}^{\ell}_{i,t})^{2}.$ (4)
This energy function is exactly the variational free energy defined in the previous section, and a proper derivation can be found in [56]. Let us assume that our generative model is presented with an observation $\mathbf{o}\in\mathbb{R}^{J_{0}}$. Then, the following process describes the credit assignment and consequent update of the synaptic parameters of the model: first, the neurons of the lowest layer are set equal to the sensory observation, i.e., $\mathbf{x}^{0}=\mathbf{o}$. Next, the unconstrained neural activities are updated until the convergence step $T$ (or, for a fixed number $T$ of iterations) to minimize the energy of Eq. 4 via gradient descent. In particular, the equation for the update dynamics of the value nodes is the following:
@@
where $\alpha$ is the learning rate for the synaptic update. The alternation of these two phases, i.e., the value node update and weight update steps, defines the learning algorithm used to train PC networks [20, 53]. Importantly, although every computation is local, both update rules minimize the same energy function, which is globally defined over the entire network, as described in 1. The difference in locality between a weight update of backprop and the one for PC is given in Fig. 3(a).
@@
[…]
@@
3:   Update $\mathbf{x}^{\ell}_{i,t}$ to minimize $\mathcal{E}_{t}$ via Eq. (5)
@@
In the above, we have described how to use PC to perform unsupervised learning tasks. However, this algorithm or variations of it have been shown to obtain a performance comparable to deep networks learned with backprop on supervised learning tasks [53, 57, 58]. To extend the above formulation to the domain of labeled data, we have to re-imagine the PC network as generating the label $\mathbf{y}$ and that the data point $\mathbf{o}$ serves as a prior on the value of the neurons in the first layer of the network. Let us assume that we have a labeled data point $(\mathbf{o},\mathbf{y})$, where, again, $\mathbf{y}$ is the label  [^1] In supervised learning, the total energy of the model can also be decomposed as $\mathcal{E}=\tilde{\mathcal{E}_{t}}$$\mathcal{L}$, with $\tilde{\mathcal{E}_{t}}$ be the total energy of all the internal neurons only, and $\mathcal{L}$ be the squared error defined on the label, as in standard regression tasks. This allows to define interesting similarities with backprop [59].. Supervised learning is then realized by fixing the value nodes of the first and last layer to the entries of the data point and its label, respectively, i.e., $\mathbf{x}_{0,t}=\mathbf{o}$ and $\mathbf{x}_{L,t}=\mathbf{y}$ for every time step $t$. A graphical depiction of the difference between supervised and unsupervised learning using PC is given in Fig. 4.
@@
While defining this generalization on fully connected graphs makes the notation simpler, interesting use cases can be found in graphs that are not fully connected. To this end, it suffices to note that every graph is a subset of a fully connected one and, hence, we can consider sparse weight matrices $\mathbf{W}$, where only the entries of the edges that we need to define our structure are non-zero. This can be done by multiplying $\mathbf{W}$ by an adjacency matrix. In the original work along these lines, the authors trained a network of multiple, sparsely connected brain regions and tested the resulting model on image denoising and image completion tasks. The exact structure of this network is that of the Assembly Calculus, a Hebbian learning framework/method specifically designed to model brain regions [61]. A graphical sketch of how to derive such a model is given in Fig. 3(b).
@@
A body of literature has investigated the efficacy of PC networks in memory-related tasks. Specifically, this literature studies their capacity to store a dataset and retrieve individual data points when a related cue is presented. In practice, this cue is often an incomplete or corrupted version of a stored memory, and the task is successfully completed if and only if the correct data point is retrieved. PC schemes have demonstrated their ability to effectively store and retrieve complex memories, such as images from the COIL and ImageNet datasets [62, 63, 64]. However, while their retrieval capabilities are robust, their capacity is still limited, and hence not comparable to that of continuous or universal Hopfield networks [65, 66]. This capacity can be improved by implementing a fast and powerful memory writing operation that allows these models to store individual memories without overwriting existing ones. Models that incorporate memory writing also tend to facilitate the implementation of a forget operation that erases individual memories with little to no impact on overall model performance [67]. It is also possible to store memories in the latent variable $\mathbf{x}_{L}$, by learning a Gaussian mixture prior, whose centers correspond to each of the individual memories [68]. This has also been implemented in the temporal domain, where the goal is to retrieve missing future frames of a video, given initial ones [64] (as a sort of priming stimulus). Note, however, that this class of models does not have the exponential capacity of modern Hopfield networks [66, 69].
3.1 Differences and Similarities with Backpropagation
In the case of supervised learning, PC can be seen simply as a training algorithm for deep neural networks that can be used as an alternative to backprop. However, the two algorithms differ in terms of convergence, complexity, and performance, in ways that we now discuss.
Similarities.
@@ 3.1 Differences and Similarities with Backpropagation
Recent work has shown that PC can approximate the weight update of backprop, under specific conditions, on both MLPs (multilayer perceptrons) and within the more general framework of computational graphs [53, 44, 70, 71]. These conditions are restrictive in practice, as these results only hold if either the total prediction error on the network is infinitesimally small or the predictions are kept constant throughout the whole duration of the inference process, i.e., $\mathbf{u}^{\ell}_{i,t}=\mathbf{u}^{\ell}_{i,0}$ for every time step $t$. In practice, however, empirical studies have shown that for the approximation to hold, it suffices to have a small output error [44]. A similar line of work has also shown that simply adding a temporal scheduling to the updates of the weights and the neural activities leads to a weight update that is equivalent to that of backprop [72, 73]. This temporal scheduling, that sequentially updates the weights of different layers at different, pre-defined, time steps, is however implausible, as it requires external control that triggers the updates at different time steps [74].
Differences.
@@
PC has been shown to perform better than standard models (such as backprop-trained deep neural networks) on problems that are faced by biological organisms, such as continual learning, online learning, and learning from a small amount of data [75, 76, 27]. This is due to the inference phase, which allows the error to be distributed in the network in a way that avoids a phenomenon called weight interference. A second difference is about stability and convergence: in backprop, and in models trained with gradient descent on a non-local loss function, each parameter update is computed independently based on the current network state, without accounting for how other parameters will simultaneously change. This has been shown to produce updates that destabilize training [77]. This is not true for PC networks, as it has been shown that PC models trained for supervised learning naturally implement implicit gradient descent [28], a more stable optimization approach where each parameter update takes into account the simultaneous changes that will occur to all other parameters in the network. Specifically, a parameter update of implicit gradient descent is defined as follows:
@@
This formulation is called implicit, since $\mathbf{W}_{t+1}$ appears on both sides of the equation, desirably reducing the sensitivity to the learning rate [78]. As a consequence, PC models tend to be more robust and better calibrated than standard ones, as has been shown in multiple classification tasks on both convolutional and graph networks [60, 79, 57]. Recent works have also studied and analyzed the convergence of PC models, showing that the iterative inference phase allows them to better escape bad local minima and saddle points, as well as needing a smaller number of weight updates with respect to backpropagation [29, 80].
Limitations.
Differently from backprop-based models, PCNs face a scalability bottleneck that limits their effectiveness on large-scale deep learning tasks, preventing any kind of industrial adoption. For example, in classification tasks, these models achieve competitive performance when tested on standard benchmarks with shallow architectures, such as convolutional models with 5-7 layers. However, we observe that performance degrades significantly as network depth increases [81]. This has been shown both in PC models trained with gradient descent, and in models trained via closed-form updates [82]. In both cases, the core issue stems from energy propagation dynamics: deeper networks experience exponentially larger energy magnitudes in layers closer to the output, creating unstable training conditions where parameter updates concentrate in a small subset of layers rather than utilizing the full network depth. This energy imbalance causes latent states to diverge excessively from forward pass activations, leading to suboptimal weight updates and poor credit assignment throughout the network. While strategies such as energy regularization, modified weight update schemes, and initialization strategies are improving the results [83, 84, 85], there is still a significant amount of research to be performed to be able to make PC match the performance of BP on larger architectures, such as language models [86].
A second limitation concerns the computational complexity of the inference phase. Theoretically, optimal performance requires waiting until convergence before updating weights, which introduces substantial computational overhead. In practice, however, a limited number of inference iterations yields competitive results, typically requiring iterations at least equal to the network depth, with empirical evidence suggesting that optimal performance occurs when $L<T<2L$ iterations are performed [81]. When the iteration count approximates the layer count, the time complexity becomes comparable to backprop, though space complexity remains higher due to the need to maintain all neurons and layers simultaneously in memory. This computational profile creates significant engineering challenges, particularly the absence of deep learning libraries capable of fully parallelizing the inference phase. Current implementations fall into two distinct categories: libraries that achieve full parallelization but lack seamless integration with established deep learning frameworks like PyTorch and Equinox, such as NGC-Learn [24] and jPC [87]; and libraries that provide modular, framework-compatible approaches but sacrifice parallelization, such as PCX [81].